OpenStudy (anonymous):

Simplify, if possible. log2(0.5)4

OpenStudy (anonymous):

Simplify, if possible. log2(0.5)^4

OpenStudy (agreene):

1

OpenStudy (agreene):

\[[\log_2(0.5)]^4=[\frac{\ln(0.5)}{ln(2)}]^4=[-1]^4=1\]

OpenStudy (anonymous):

i figured it out. it was -4

OpenStudy (hoblos):

yes it is -4 log2(0.5)^4 = 4log2(1/2) = 4[log2(1) - log2(2)] = 4(-1)=-4